OpenStudy (sonic):

How do you show that i*i=j*j=k*k=1 for vectors/dot product?

OpenStudy (anonymous):

i is a unit vector so l i l =1 i * i= l i l * l i lcos(0) =1*1*1=1 similarly j*J=k*k=1

OpenStudy (sonic):

Is i being a unit vector a basic definition or based on another proof?

OpenStudy (anonymous):

its a basic definition

OpenStudy (sonic):

Got it, thanks!